# MOLLM: Multi-Objective Large Language Model for Molecular Design – Optimizing with Experts

### v1

@@ Abstract
Molecular design plays a critical role in advancing fields such as drug discovery, materials science, and chemical engineering. This work introduces the Multi-Objective Large Language Model for Molecular Design (MOLLM), a novel framework that combines domain-specific knowledge with the adaptability of Large Language Models to optimize molecular properties across multiple objectives. Leveraging in-context learning and multi-objective optimization, MOLLM achieves superior efficiency, innovation, and performance, significantly surpassing state-of-the-art (SOTA) methods. Recognizing the substantial impact of initial populations on evolutionary algorithms, we categorize them into three types: best initial, worst initial, and random initial, to ensure the initial molecules are the same for each method across experiments. Our results demonstrate that MOLLM consistently outperforms SOTA models in all of our experiments. We also provide extensive ablation studies to evaluate the superiority of our components.
@@ 1 Introduction
Molecular design is fundamental in fields such as drug discovery, materials science, and chemical engineering. In these areas, the ability to design novel molecules with targeted properties, including stability, reactivity, or bioactivity, can drive significant advancements, from the development of new pharmaceuticals to the creation of sustainable, innovative materials. Traditionally, molecular design has relied on trial-and-error experimentation and repeated synthesis, which is resource intensive, time-consuming, and ultimately inefficient. During the past few decades, with rapid advances in computational power, various machine learning techniques (Elton et al., 2019; Du et al., 2022) have been introduced not only to accelerate this process but also to enable the discovery of novel, more potent molecules. Methods include Bayesian Optimization (BO) (Tripp et al., 2021), Multi-Objective Optimization (MOO) (Liu et al., ; Choi et al., 2023; Verhellen, 2022), Markov Chain Monte Carlo (MCMC) (Xie et al., 2021; Sun et al., 2022), Genetic Algorithms (GA) (Jensen, 2019; Nigam et al., 2019; Liu et al., 2024b; Brahmachary et al., 2024; Wang et al., 2024b, a), Reinforcement Learning (RL) (Olivecrona et al., 2017; Jin et al., 2020; Fu et al., 2022), and Deep Learning (DL) models (Jin et al., 2018b; Bagal et al., 2021; Lee et al., 2023; Fang et al., 2024).
@@
Although these methods have yielded excellent results, most of them lack the integration of expert knowledge during runtime, despite the crucial role of professional feedback and search direction in molecular design. Large Language Models (LLMs), typically based on transformer architectures (Vaswani, 2017), are pre-trained on extensive high-quality data, including books and academic papers, enabling them to capture domain-specific expertise. They have demonstrated significant potential in scientific discovery, particularly in molecular understanding and the generation of novel molecular candidates, as exemplified by models like GPT-4 (AI4Science & Quantum, 2023). Recent studies highlight the advantages of in-context learning (Nguyen & Grover, 2024) and iterative evolutionary approaches (Wang et al., 2024b) in enhancing LLM effectiveness. However, research in this area remains nascent, with only preliminary findings and a lack of systematic investigation.
@@
Furthermore, despite significant progress in training large neural networks to understand chemistry and molecular structures with domain knowledge, these models often require additional parameters and retraining, particularly for MOO, as seen in MolGPT (Bagal et al., 2021) and LICO (Nguyen & Grover, 2024). In contrast, MOLLEO (Wang et al., 2024b) leverages domain knowledge from pre-trained large language models without additional training but still relies on GB-GA within its framework.
In practice, most molecular design tasks optimizes multiple objectives, yet existing methods often ignore this aspect. For example, GB-BO (Tripp et al., 2021), JTVAE (Jin et al., 2018a), and MolGen (Fang et al., 2024) provide limited multi-objective capabilities. MolGPT requires specific training for different objectives, restricting its flexibility, while MolGen focuses primarily on target molecular discovery and employs only single-objective optimization.
Finally, the formulation of MOO using GA has often lacked rigor in previous studies. First, oracle calls should be restricted to ensure fair comparisons and practical applicability, since the evaluation of certain molecular properties requires costly experiments or specifically trained models, as noted in the Practical Molecular Optimization benchmark (Gao et al., 2022). Additionally, the initial population significantly impacts final performance under a fixed number of oracle calls, yet this factor has been largely overlooked in methods that incorporate genetic algorithms, such as MARS, MOLLEO, and GB-GA.
@@
We carefully design the in-context learning and prompt engineering mechanism in our model to fully leverage the domain knowledge pre-trained in LLMs. This is seamlessly integrated into MOO framework, achieving SOTA performance in both optimization quality and efficiency. Our framework requires no additional training for specific objectives while capitalizing on domain expertise, reasoning capabilities, and is adaptable to various LLMs. Unlike MOLLEO, we employ LLMs for all mating operations, ensuring that the framework is entirely LLM-driven.
Recognizing the critical influence of initial population selection in genetic algorithm-based methods, we evaluate models using three types of initial populations: the worst, random, and best molecules from the ZINC250K dataset. Our results show that MOLLM outperforms all kinds of SOTA models built on GA, BO, MCMC, LLM, RL and DL in our experiments, particularly in maximizing the sum of absolute property values in multi-objective settings. In addition, extensive ablation studies validate the effectiveness of our approach and design choices.
@@ 2.1 Molecular Design with Machine Learning
Numerous advanced models for molecular design span GA, BO, MOO, MCMC, RL, and DL methodologies. Deep Learning (DL) leverages neural networks in various molecular design models. Differentiable Scaffolding Tree (DST) (Fu et al., 2021) with GNNs, Junction Tree Variational Autoencoders (JTVAE) (Jin et al., 2018a), and VJTNN+GAN (Jin et al., 2018b) combine generative and adversarial architectures to generate molecules. MOOD (Lee et al., 2023) utilizes Diffusion models to address out-of-distribution generation. Recent developments in Generative Pre-trained Transformers (GPT) led Bagal et al. to train MolGPT (Bagal et al., 2021) on next-token prediction, while Fang et al. pre-trained MOLGEN (Fang et al., 2024) on molecule reconstruction tasks, achieving cross-domain applicability. Although DL methods offer powerful capabilities in capturing complex molecular structures and enabling cross-domain applicability such as DST, JTVAE, and MolGPT, they often underperform in MOO scenario. Latent Space Optimization (LSO) (Abeer et al., 2024) has further advanced multi-objective molecular design, but only for deep generative models.
@@
Reinforcement Learning (RL) combined with DL iteratively refines molecules by learning from feedback, often based on property scores. REINVENT (Olivecrona et al., 2017) applies RL to train an RNN to generate molecules meeting multiple goals, while RationaleRL (Jin et al., 2020) uses a Graph Neural Network (GNN) to generate molecules by building interpretable substructures, or “rationales”. Based on REINVENT, Shin et al. proposed a novel divide-and-conquer approach called DyMol (Shin et al., 2024) to train the model for multiple objectives and achieve SOTA results. Kim et al. also achieve SOTA performance by integrating genetic algorithms into GFlowNets (Kim et al., 2024).
@@
In addition to DP methods, classical probabilistic models and optimization methods also achieve SOTA performance in many cases, such as in PMO (Gao et al., 2022). A notable example of Genetic Algorithms (GA) is GB-GA (Jensen, 2019), commonly used as a baseline, where molecular structures are modified in graph form during mating operation. AkshatKumar et al. (Nigam et al., 2019) introduced a neural network discriminator to enhance diversity, surpassing GB-GA in maximizing penalized-logP (Gómez-Bombarelli et al., 2018). Later, Tripp et al. (Tripp et al., 2021) employed a Tanimoto kernel in a Gaussian Process in GB-GA, outperforming GB-GA. It uses SELFIES (Krenn et al., 2020), a 100% valid molecular representation system; however, Gao et al. (Gao et al., 2022) later showed there are no obvious shortcomings of SMILES compared to SELFIES. MLPS (Liu et al., ) combines MOO with BO and an encoder-decoder network to efficiently locate global Pareto-optimal solutions, while Verhellen et al. introduced a graph-based MOO (Verhellen, 2022) for molecular optimization. Furthermore, MARS (Xie et al., 2021) uses Markov Chain Monte Carlo (MCMC) to explore chemical spaces probabilistically to identify molecules with desirable properties. Similarly, MolSearch (Sun et al., 2022) utilizes Monte Carlo tree search for multi-objective molecular discovery. However, GA, BO, MOO, and MCMC methods are independent of domain knowledge, which is highly beneficial in molecular design but challenging to incorporate into such algorithms.
@@ 2.2 Multi-Objective Optimization and Genetic Algorithm with LLM
Recently, Large Language Models (LLMs) have demonstrated remarkable performance across various Natural Language Processing (NLP) benchmarks (Brown, 2020; AI4Science & Quantum, 2023), sparking interest in their application as optimization operators in MOO to address the challenges of high-dimensional search spaces and to incorporate domain knowledge (Wu et al., 2024). For instance, OPRO (Yang et al., 2024) and LMEA (Liu et al., 2024b) employ LLMs as crossover and mutation operators within GA, using prompts that include parent values from the current population, with LMEA further exploring the balance of exploitation and exploration by adjusting the temperature parameter. Furthermore, Wang et al. (Wang et al., 2024c) investigated constrained MOO with prompt engineering, demonstrating promising alignment results. Other studies have highlighted the effectiveness and efficiency of LLMs in GA compared to standalone LLMs and traditional MOO algorithms, proposing well-structured pipelines (Liu et al., 2023, 2024a, 2024c; Huang et al., 2024; Brahmachary et al., 2024). However, research on LLMs with MOO is still nascent, with results largely limited to preliminary findings in numerical optimizations and planning problems.
@@ 2.3 Molecular Design with LLM
LLMs with pre-trained domain knowledge are increasingly popular for accelerating drug discovery and materials design (AI4Science & Quantum, 2023). In particular, ChemCrow (M. Bran et al., 2024) uses LLMs as agents capable of reasoning, planning, and selecting appropriate external tools to iteratively refine candidates in chemical tasks. LICO (Nguyen & Grover, 2024) improves molecule generation through in-context learning by pretraining the model with separate embedding and prediction layers, while Moayedpour et al. (Moayedpour et al., 2024) extend this approach to multi-objective setups, and MolReGPT (Li et al., 2024) advances few-shot learning for molecular optimization. MOLLEO (Wang et al., 2024b) applies GA combined with LLMs for molecular design, aligning with our framework, but differing significantly in details. MOLLEO’s results as well as research in prompts remain preliminary, lacking extensive multi-objective experiments and failing to consider the impact of varying initial populations.
@@ 3.1 MOLLM Overview
Figure 1 presents the complete MOLLM optimization pipeline. The task involves unconstrained molecular optimization, where, given a set of objectives, the model is initialized with molecules selected from the ZINC250K dataset. It then iterates through mating, prompt generation, scoring, experience updating, and next-generation selection. Initialization: For an optimization problem with one or multiple objectives, we initialize with molecules, either randomly selected or chosen based on the best or worst objective values from the ZINC250K dataset. In our setting, . The ZINC dataset (Irwin et al., 2012) is selected as it is widely used for population initialization in molecular optimization studies (Wang et al., 2024b; Jensen, 2019; Fu et al., 2022). ZINC250K comprises approximately 250,000 curated drug-like molecules from the ZINC database, providing key properties such as chemical structures, logP, QED, and SA, making it well-suited for drug discovery and molecular optimization tasks.
Mating: This step involves prompting the LLM to generate new candidate molecules that are expected to improve on the parent molecules given. First, the parent molecules are randomly selected from the current population with probabilities and for crossover and mutation, respectively. Each crossover involves two parents, while mutation involves a single parent. These selected parents are then formatted into a flexible prompt template. Our prompt template consists of five key components: multi-objective requirements, objective descriptions, parent objective values (for in-context learning), output instructions, and past experience, if applicable. The whole framework of our model and an example is shown in Figure 1. In this setup, the model receives structured input specifying primary objectives, descriptions of molecular modifications (in SMILES format) that may increase or decrease property values, and parent molecules represented by their SMILES structures, objective values, and an aggregated objective score as an overall performance indicator. The output instructions specify that only molecular structures should be generated, omitting explanations to significantly reduce runtime and query costs without affecting performance. We employ crossover and mutation to balance exploitation and exploration. Although LLMs perform well in crossover due to their straightforward nature, they struggle with mutation, as its prompt is highly similar to crossover. To address this, we provide a list of common molecular mutation operations in the instructions to improve exploration. Finally, after generating offspring, we identify the best and worst molecules among them and query the LLM to update its experience based on these molecules and experience in the last iteration. This iterative refinement allows the experience to evolve dynamically, transitioning from general suggestions to more detailed and actionable guidance over time.

### v2

@@ Abstract
Molecular design plays a critical role in advancing fields such as drug discovery, materials science, and chemical engineering. This work introduces the Multi-Objective Large Language Model for Molecular Design (MOLLM), a novel framework that combines domain-specific knowledge with the adaptability of large language models to optimize molecular properties across multiple objectives. Leveraging in-context learning and multi-objective optimization, MOLLM achieves superior performance and innovation, consistently surpassing state-of-the-art (SOTA) methods. We significantly improve the efficiency of our framework, making it 14 times faster and substantially more cost-effective without compromising performance compared to the latest similar work. Our results demonstrate that MOLLM consistently outperforms SOTA models across experiments and excels on the PMO benchmark. In addition, we provide extensive ablation studies and analysis to evaluate the effectiveness of each component and the quality of the output molecules.
@@ 1 Introduction
Molecular design is fundamental in fields such as drug discovery, materials science, and chemical engineering. In these areas, the ability to design novel molecules with targeted properties, including stability, reactivity, or bioactivity, can drive significant advancements, from the development of new pharmaceuticals to the creation of sustainable, innovative materials. Traditionally, molecular design has relied on trial-and-error experimentation and repeated synthesis, which is resource intensive, time-consuming, and ultimately inefficient. During the past few decades, with rapid advances in computational power, various machine learning techniques Elton et al. (2019); Du et al. (2022) have been introduced not only to accelerate this process but also to enable the discovery of novel, more potent molecules. Methods include Bayesian Optimization (BO) Tripp et al. (2021), Multi-Objective Optimization (MOO) Liu et al. ; Choi et al. (2023); Verhellen (2022), Markov Chain Monte Carlo (MCMC) Xie et al. (2021); Sun et al. (2022), Genetic Algorithms (GA) Jensen (2019); Nigam et al. (2019); Liu et al. (2024b); Brahmachary et al. (2024); Wang et al. (2024b, a), Reinforcement Learning (RL) Olivecrona et al. (2017); Jin et al. (2020); Fu et al. (2022), and Deep Learning (DL) models Jin et al. (2018b); Bagal et al. (2021); Lee et al. (2023); Fang et al. (2024).
@@
Although these methods have yielded excellent results, most of them lack the integration of expert knowledge during runtime, despite the crucial role of professional feedback and search direction in molecular design. Large Language Models (LLMs), typically based on transformer architectures Vaswani (2017), are pre-trained on extensive high-quality data, including books and academic papers, enabling them to capture domain-specific expertise. They have demonstrated significant potential in scientific discovery, particularly in molecular understanding and the generation of novel molecular candidates, as exemplified by models like GPT-4 AI4Science and Quantum (2023). Recent studies highlight the advantages of in-context learning Nguyen and Grover (2024) and iterative evolutionary approaches Wang et al. (2024b) in enhancing LLM effectiveness. However, research in this area remains nascent, with only preliminary findings and a lack of systematic investigation.
@@
Furthermore, despite significant progress in training large neural networks to understand chemistry and molecular structures with domain knowledge, these models often require additional parameters and retraining, particularly for MOO, as seen in MolGPT Bagal et al. (2021) and LICO Nguyen and Grover (2024). In contrast, MOLLEO Wang et al. (2024b) leverages domain knowledge from pre-trained large language models without additional training but still relies on GB-GA within its framework and only uses a weighted sum for solving multi-objective tasks. In practice, most molecular design tasks optimizes multiple objectives, yet existing methods often ignore this aspect. For example, GB-BO Tripp et al. (2021), JTVAE Jin et al. (2018a), and MolGen Fang et al. (2024) provide limited multi-objective capabilities. MolGPT requires specific training for different objectives, restricting its flexibility, while MolGen focuses primarily on target molecular discovery and employs only single-objective optimization.
@@
We carefully design the in-context learning and prompt engineering mechanisms to fully leverage the domain knowledge encoded in LLMs. These components are seamlessly integrated into the MOO framework, achieving SOTA performance on the PMO benchmark and additional multi-objective optimization tasks in terms of both quality and efficiency. Our framework requires no task-specific training, benefits from domain reasoning capabilities, and is adaptable to various LLMs. Unlike MOLLEO, our model employs LLMs for all mating operations, making the entire framework LLM-driven. This model is easy to use and can effectively accelerate the process of discovering novel drugs.
We develop an experience pool that is refined iteratively during optimization. This memory mechanism accelerates convergence and improves optimization results by enhancing the model’s exploitation ability. Moreover, it improves the explainability of the optimization process and provides reusable and insightful experience at the end of each experiment.
We introduce several modifications to enhance exploration and balance it with exploitation: (1) We identify that the number of output molecules clearly influences the LLM’s exploratory capacity and determine an optimal output number (illustrated in Section 7.3 in Appendix). (2) We employ Pareto front selection instead of only using simple objective summation, which promotes diversity in search directions. These components contribute to increased performance as the number of objectives grows compared to MOLLEO. By further reducing redundant components in the framework, we achieve approximately 14× faster than MOLLEO and requires significantly fewer LLM queries, without compromising performance.
@@ 2.1 Molecular Design with Machine Learning
Numerous advanced models for molecular design span GA, BO, MOO, MCMC, RL, and DL methodologies. Deep Learning (DL) leverages neural networks in various molecular design models. Differentiable Scaffolding Tree (DST) Fu et al. (2021) with GNNs, Junction Tree Variational Autoencoders (JTVAE) Jin et al. (2018a), and VJTNN+GAN Jin et al. (2018b) combine generative and adversarial architectures to generate molecules. MOOD Lee et al. (2023) utilizes Diffusion models to address out-of-distribution generation. Recent developments in Generative Pre-trained Transformers (GPT) led Bagal et al. to train MolGPT Bagal et al. (2021) on next-token prediction, while Fang et al. pre-trained MOLGEN Fang et al. (2024) on molecule reconstruction tasks, achieving cross-domain applicability. Although DL methods offer powerful capabilities in capturing complex molecular structures and enabling cross-domain applicability such as DST, JTVAE, and MolGPT, they often underperform in MOO scenario. Latent Space Optimization (LSO) Abeer et al. (2024) has further advanced multi-objective molecular design, but only for deep generative models.
@@
Reinforcement Learning (RL) combined with DL iteratively refines molecules by learning from feedback, often based on property scores. REINVENT Olivecrona et al. (2017) applies RL to train an RNN to generate molecules meeting multiple goals, while RationaleRL Jin et al. (2020) uses a Graph Neural Network (GNN) to generate molecules by building interpretable substructures, or “rationales”. Based on REINVENT, Shin et al. proposed a novel divide-and-conquer approach called DyMol Shin et al. (2024) to train the model for multiple objectives and achieve SOTA results. Kim et al. also achieve SOTA performance by integrating genetic algorithms into GFlowNets Kim et al. (2024). Guo and Schwaller achieve SOTA performance by introducing Augmented Memory, which integrates experience replay and SMILES augmentation for efficient molecular optimization Guo and Schwaller (2024).
@@
In addition to DP methods, classical probabilistic models and optimization methods also achieve SOTA performance in many cases, such as in PMO Gao et al. (2022). A notable example of Genetic Algorithms (GA) is GB-GA Jensen (2019), commonly used as a baseline, where molecular structures are modified in graph form during mating operation. AkshatKumar et al. Nigam et al. (2019) introduced a neural network discriminator to enhance diversity, surpassing GB-GA in maximizing penalized-logP Gómez-Bombarelli et al. (2018). Later, Tripp et al. Tripp et al. (2021) employed a Tanimoto kernel in a Gaussian Process in GB-GA, outperforming GB-GA. It uses SELFIES Krenn et al. (2020), a 100% valid molecular representation system; however, Gao et al. Gao et al. (2022) later showed there are no obvious shortcomings of SMILES compared to SELFIES. MLPS Liu et al. combines MOO with BO and an encoder-decoder network to efficiently locate global Pareto-optimal solutions, while Verhellen et al. introduced a graph-based MOO Verhellen (2022) for molecular optimization. Furthermore, MARS Xie et al. (2021) uses Markov Chain Monte Carlo (MCMC) to explore chemical spaces probabilistically to identify molecules with desirable properties. Similarly, MolSearch Sun et al. (2022) utilizes Monte Carlo tree search for multi-objective molecular discovery. However, GA, BO, MOO, and MCMC methods are independent of domain knowledge, which is highly beneficial in molecular design but challenging to incorporate into such algorithms.
@@ 2.2 Multi-Objective Optimization and Genetic Algorithm with LLM
Recently, Large Language Models (LLMs) have demonstrated remarkable performance across various Natural Language Processing (NLP) benchmarks Brown (2020); AI4Science and Quantum (2023), sparking interest in their application as optimization operators in MOO to address the challenges of high-dimensional search spaces and to incorporate domain knowledge Wu et al. (2024a). For instance, OPRO Yang et al. (2024) and LMEA Liu et al. (2024b) employ LLMs as crossover and mutation operators within GA, using prompts that include parent values from the current population, with LMEA further exploring the balance of exploitation and exploration by adjusting the temperature parameter. Furthermore, Wang et al. Wang et al. (2024c) investigated constrained MOO with prompt engineering, demonstrating promising alignment results. Other studies have highlighted the effectiveness and efficiency of LLMs in GA compared to standalone LLMs and traditional MOO algorithms, proposing well-structured pipelines Liu et al. (2023, 2024a, 2024c); Huang et al. (2024); Brahmachary et al. (2024). However, research on LLMs with MOO is still nascent, with results largely limited to preliminary findings in numerical optimizations and planning problems.
@@ 2.3 Molecular Design with LLM
LLMs with pre-trained domain knowledge are increasingly popular for accelerating drug discovery and materials design AI4Science and Quantum (2023). In particular, ChemCrow M. Bran et al. (2024) uses LLMs as agents capable of reasoning, planning, and selecting appropriate external tools to iteratively refine candidates in chemical tasks. LICO Nguyen and Grover (2024) improves molecule generation through in-context learning by pretraining the model with separate embedding and prediction layers, while Moayedpour et al. Moayedpour et al. (2024) extend this approach to multi-objective setups, and MolReGPT Li et al. (2024) advances few-shot learning for molecular optimization. MOLLEO Wang et al. (2024b) applies GA combined with LLMs for molecular design, aligning with our framework, but differing significantly in details. MOLLEO’s results as well as research in prompts remain preliminary, lacking specific consideration and modification for multi-objective settings. Prompt-MolOpt Wu et al. (2024b) explicitly leverages the domain knowledge of LLMs for molecular design and demonstrates strong zero- and few-shot capabilities through property-specific prompt embeddings combined with a sequence-to-sequence Transformer trained on substructure-annotated molecule pairs. However, its practical evaluations are mostly limited to two-objective settings, and its performance on more complex combinations still depends on property-specific prompt engineering and single-objective training data.
@@ 3.1 MOLLM Overview
Figure 4 presents the complete MOLLM optimization pipeline. The task involves unconstrained molecular optimization, where, given a set of objectives, the model is initialized with molecules selected from the ZINC250K dataset. It then iterates through mating, prompt generation, scoring, experience updating, and next-generation selection. Initialization: For an optimization problem with one or multiple objectives, we initialize with molecules, either randomly selected or chosen based on the best or worst objective values from the ZINC250K dataset. In our setting, . The ZINC dataset Irwin et al. (2012) is selected as it is widely used for population initialization in molecular optimization studies Wang et al. (2024b); Jensen (2019); Fu et al. (2022). ZINC250K comprises approximately 250,000 curated drug-like molecules from the ZINC database, providing key properties such as chemical structures, logP, QED, and SA, making it well-suited for drug discovery and molecular optimization tasks.
Mating: This step involves prompting the LLM to generate new candidate molecules that are expected to improve on the parent molecules given. First, the parent molecules are randomly selected from the current population with probabilities and for crossover and mutation, respectively. Each crossover involves two parents, while mutation involves a single parent. These selected parents are then formatted into a flexible prompt template. Our prompt template consists of five key components: multi-objective requirements, objective descriptions, parent objective values (for in-context learning), output instructions, and past experience. The whole framework of our model and an example is shown in Figure 4. In this setup, the model receives structured input specifying primary objectives, descriptions of molecular modifications (in SMILES format) that may increase or decrease property values, and parent molecules represented by their SMILES structures, objective values, and an aggregated objective score as an overall performance indicator. The output instructions specify that only molecular structures should be generated, omitting explanations to significantly reduce runtime and query costs without affecting performance. We employ crossover and mutation to balance exploitation and exploration. Although LLMs perform well in crossover due to their straightforward nature, they struggle with mutation, as its prompt is highly similar to crossover. To address this, we provide a list of common molecular mutation operations in the instructions to improve exploration. Finally, after generating offspring, we identify the best and worst molecules among them and query the LLM to update its experience based on these molecules and experience in the last iteration. This iterative refinement allows the experience to evolve dynamically, transitioning from general suggestions to more detailed and actionable guidance over time.